Circular No. 7035

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1998 U2 J. Mueller reports her discovery of another comet on plates taken by

herself (and K. Rykoski on Oct. 22) with the 1.2-m Oschin Schmidt Telescope in the course of the Palomar Outer Solar System Ecliptic Survey. The comet has strong condensation and a short, faint tail to the southeast. The following positions were reduced by G. V. Williams, Smithsonian Astrophysical Observatory, from x, y measures provided by Mueller:

1998 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 21.34028  1^{h}24^{m}56.^{s}57  +12^{o}27′57′′.2  18
21.36111  1 24 56.12  +12 27 55.1
22.20417  1 24 30.07  +12 26 17.8
22.22500  1 24 29.65  +12 26 15.2

SUPERNOVAE 1998ej AND 1998ek IN ANONYMOUS GALAXIES J. Maza, Department of Astronomy, University of Chile, reports the

discovery of two supernovae by Roberto Antezana on T-Max 400 films taken by L. E. González with the Maksutov telescope at Cerro El Roble:

SN  1998 UT  α_{2000}  δ_{2000}  Mag.  Offset
1998ej Oct. 15.201 23^{h}41^{m}00^{s}.95 −38^{o}56′37′′.2 19  4′′.1 W, 6′′.5 N
1998ek Oct. 15.341 6 36 30.59 −37 58 40.2 18 14′′.0 E, 4′′.3 N

SN 1998ej was confirmed via B and V direct images taken by D. González on Oct. 17 with the YALO telescope at Cerro Tololo, but it is not visible on a film obtained at Cerro El Roble on Sept. 28.223 (B > 19.5). SN 1998ek was confirmed on a T-Max 400 film taken by L. E. González on Sept. 28.378 with the Cerro El Roble telescope, but it is not visible on a film obtained with the same instrument on Aug. 28.379 (B > 19.5).

COMET C/1998 T1 (LINEAR) Total visual magnitude estimates: Oct. 13.77 UT, 14.6 (M. Lehký,

Hradec Králové, Czech Republic, 0.42-m reflector); 13.85, 14.3 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 17.03, 14.5 (Lehký); 20.90, 13.7 (Lehký); 21.83, 13.9 (Hornoch).

1998 October 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green